# Uncertainty-Aware Search and Value Models: Mitigating Search Scaling Flaws in LLMs

### v1

@@ 5.2 Implementation
We train the base models (i.e. Mistral 7B) on the official training set. We use the newline character as the marker for the end of each step. Supervised fine-tuning is performed for 2 epochs with a batch size of 128. We use a linear learning rate scheduler with a maximum learning rate of 2e-6. The AdamW optimizer Loshchilov and Hutter (2019) is used for training.
@@
The dataset construction process is introduced in Appendix A.1. We sample 50 solution paths per problem in the training set. We use a decoding temperature of 0.7 and top-k set to 50 for dataset collection. The maximum new token length is 400. We apply vllm Kwon et al. (2023) to accelerate the generation process.
@@
To construct UVMs, we set the number of components to 10. UVMs are initialized from the corresponding generator checkpoints and trained for one epoch, using the same backbone learning rate scheduler. The maximum learning rate for the uncertainty-aware value head is set to 2e-3, with a batch size of 128. The optimizer used for training is AdamW. After training, we derive OVMs by setting $\boldsymbol{\zeta}=\mathbf{0}$.
@@
The hyperparameters for decoding are consistent with those used in the UVMs’ training dataset collection. The maximum number of steps is 10.
@@ 5.3 Overall Performance
We present the overall performance in Figure 1, showing that UVM-guided search consistently outperforms conventional OVM-guided search both when scaling sample sizes and scaling candidate sizes.
Scaling flaws
UVM-guided search mitigates the scaling flaws of conventional VM-guided search. As shown in Figure 1(a), VM-guided search leads to over 20% higher coverage than repeated sampling when the sample size is 1. However, its advantage diminishes as the sample size scales to 8. In contrast, our UVM-guided search continues to outperform repeated sampling at this sample size and remains comparable even at a sample size of 16.
Moreover, as shown in Figure 1(b), the performance of conventional VM-guided search deteriorates significantly as the candidate size scales. In contrast, UVM-guided search efficiently mitigates this issue, resulting in a 5.5% performance improvement over VM-guided search at the largest tested candidate size, 256.
Convergent points of coverage
UVM-guided search ultimately converges to higher coverage than conventional VM-guided search. As illustrated in Figure 1(a), UVM-guided search outperforms conventional VM-guided search once the sample size increases to 4. It reaches the performance comparable to the best result of conventional VM-guided search (among all tested sample sizes) using only 25% of its budget. Furthermore, it ultimately achieves 4.2% higher coverage at the largest tested sample size, 32.
These results highlight the effectiveness of our method in enhancing search scaling.
@@ 5.4 Ablation Study
In this section, we conduct ablation studies on the design choices of uncertainty-aware selection and demonstrate the superiority of the Group Thompson Sampling algorithm in Figure 3.
@@
Top-1 probability is better for candidate evaluation. As shown in Figure 3, UCB ranking performs poorly, even underperforming the conventional OVM baseline. In contrast, the top-1 probability ranking surpasses the UCB, highlighting the effectiveness of using top-1 probability for candidate evaluation.
@@
Group Thompson Sampling is an effective algorithm for uncertainty-aware selection. As illustrated in Figure 3, our proposed Group Thompson Sampling algorithm achieves higher coverage than the naive top-1 probability ranking. Furthermore, by sampling candidates according to the underlying top-1 probability distribution, Group Thompson Sampling eliminates the need for explicit top-1 probability estimation, making it more efficient.
These results show the superiority of our Group Thompson Sampling algorithm: It achieves significantly higher coverage than the other baselines while requiring fewer computations.
Figure 3: Ablation on design choices of uncertainty-aware selection
6 Analysis
In this section, we conduct in-depth analyses of VM selection failures, which are the underlying causes of search scaling flaws Yu et al. (2025). Specifically, we investigate the effectiveness of UVM in mitigating selection failures across various sparsity of correct paths among candidates.
Selection stages
We define a selection stage as comprising three key components: (1) the candidate set (2) the evaluation scores assigned to each candidate (3) the selected candidates from the set.
The presence of correct paths within the candidate set depends on the quality of candidates generated during the generation stage. In this section, we focus on the effectiveness of the selection stages and exclude issues related to the generation stages. Therefore, we only consider those selection stages where correct paths are present within the candidate set.
Selection failures
When correct paths are present in the candidate set, a selection failure is identified if none of the correct paths are selected.
In this section, we analyze the distribution of selection failures and compare the effectiveness of UVM and OVM across various scenarios.
Experimental Setup
We extract all selection stages during the OVM-guided search with $b=8,K=64$, as this setup begins to experience scaling issues while retaining a reasonable computational cost for correct path labeling and target selection stage identification.
Correct path labeling To identify the correct candidate paths, we complete each partial path by rolling out 4 samples and checking whether any of them reach the correct answer. A candidate is deemed a correct path if at least one of its rollouts leads to the correct final answer.
Target selection stage identification Given the correctness of all candidate paths across all selection stages, we can filter all the selection stages where correct paths are available among the candidate set for the subsequent studies.
Analysis on Correct Path Sparsity
Correct path sparsity refers to the fraction of correct paths within the candidate set. As sparsity increases—i.e., fewer correct paths are present among the candidates—identifying the correct paths becomes more challenging for VMs Yu et al. (2025).
Similar to the analysis in Yu et al. (2025), we categorize the correct path sparsity of all targeted selection stages (where correct paths are available) into three uniform groups. We then apply both OVM and UVM selection to the candidate sets within these stages. Then we plot the distribution of selection failures, where no correct path is identified, across the three sparsity groups in Figure 4.
UVM is effective across all correct path sparsity groups. As shown in Figure 4, UVM consistently reduces the frequency of selection failures across all sparsity groups. Specifically, it reduces selection failures by 6% in the low-sparsity group, 6% in the medium-sparsity group, and 14% in the high-sparsity group.
UVM is most effective in high-sparsity regimes. Notably, although higher correct path sparsity increases the difficulty of the selection task, UVM continues to perform well, and even offers greater benefits in reducing selection failures in high-sparsity regimes (14% failure reduction). This demonstrates its ability to retain high-uncertainty yet potentially correct paths that conventional models would typically discard prematurely.
Figure 4: Comparison of failures frequency between OVM selection and UVM selection, across various correct path sparsity
@@ 7 Discussion
Quantifying uncertainty in current search-based methods has several practical implications for LLM reasoning tasks as follows: (1) Discovering correct solutions: Correct solutions may lie along paths that are underrepresented in the training data. By incorporating uncertainty, our approach increases the likelihood of discovering these solutions. (2) Improved performance with limited resources: Uncertainty-aware search methods enhance performance without a significant increase in computational resources. This is because the process of quantifying uncertainty is computationally inexpensive. (3) Adaptability to real-world applications: Real-world applications often encounter out-of-distribution data. Methods that account for uncertainty are better equipped to handle such cases, enabling more reliable performance in deployment scenarios where the data may differ from training distributions.
@@ 8 Conclusion
VM-guided search suffers from search scaling flaws due to the use of imperfect VMs, which could produce unreliable predictions when the evaluated data is underrepresented in the training data. To address these issues, we propose an uncertainty-aware search framework, including training uncertainty-aware VMs and applying uncertainty-aware selection during the search. Experiment results in GSM8K show the effectiveness of our methods.
Limitation
While uncertainty can indicate the reliability of VM predictions, it alone cannot guarantee correct predictions. This highlights that, although uncertainty can be useful, effective selection still heavily depends on the performance of the VMs. In scenarios where VMs perform poorly, such as in the more challenging MATH dataset Hendrycks et al. (2021), uncertainty alone offers limited assistance and proves ineffective. In such scenarios, improving the quality of the VMs should be prioritized over merely equipping them with uncertainty qualifications. Besides, uncertainty does not fully capture the accuracy of value predictions. It is possible for a low-uncertainty prediction to still be incorrect, a scenario that is not addressed in this paper.
@@ Appendix A Appendix
$\boldsymbol{\theta}$ Parameters of the value model backbone
@@
$\boldsymbol{\zeta}$ Index vector of dimension $m$
$p_{\boldsymbol{\zeta}}$ Index distribution
@@ A.2 Detailed UVM Structure
Given the representation $\mathbf{x}$, the sampled index $\boldsymbol{\zeta}$ is mapped to the posterior value sample as:
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(u\mathbf{W}+p_{0}\mathbf{W}_{0})\boldsymbol{\zeta}^{T}}_{\text{uncertainty estimator}}$ (8)

### v2

@@ 5.2 Implementation
We train the three base models (i.e. Mistral 7B, Qwen2.5-Math 7B, and Qwen3 8B) on the official training set of GSM8K and MATH, respectively. We use the newline character as the marker for the end of each step. Supervised fine-tuning is performed for 2 epochs with a batch size of 128. We use a linear learning rate scheduler with a maximum learning rate of 2e-6 for Mistral 7B, 2e-5 for Qwen2.5-Math 7B, and 1e-5 for Qwen3 8B. The AdamW optimizer (Loshchilov & Hutter, 2019) is used for training.
@@
The dataset construction process is introduced in Appendix A.1. We sample 50 solution paths per problem in the training set. For dataset collection, we use a decoding temperature of 0.7, top-k set to 50, and the maximum new token length set to 400 on GSM8K; we use a temperature of 1, top-p set to 0.98, and the maximum new token length set to 2000 on MATH. We apply vllm (Kwon et al., 2023) to accelerate the generation process.
@@
To construct UVMs, we set the number of components to 10. UVMs are initialized from the corresponding generator checkpoints and trained for one epoch, using the same backbone learning rate scheduler. The maximum learning rate for the uncertainty-aware value head is set to 2e-3 for Mistral 7B, 2e-4 for Qwen2.5-Math 7B, and 1e-3 for Qwen3-8B. The batch size is set to 128 on GSM8K and to 512 on MATH. The optimizer used for training is AdamW. After training, we derive OVMs by setting $\bm{\zeta}=\mathbf{0}$.
@@
The decoding hyperparameters mirror those used for UVM training data collection; we set the maximum number of steps to 10 on GSM8K and 20 on MATH, and for AIME25 and Minerva Math, we adopt the MATH settings with two exceptions: a 4096 maximum new token length and a 50 maximum number of steps.
@@ 5.3 Overall Performance
We present the overall performance in Table 1 and Table 2.
Coverage gains with UVM
UVM-guided search consistently outperforms conventional OVM-guided search on coverage in both ID and OOD settings. In ID settings, the average coverage across models and benchmarks increases by 2.2%. For example, it raises Mistral’s coverage on GSM8K by about 4.3% and Qwen2.5-Math’s coverage on MATH by about 1.3%. Gains are larger in the OOD settings: average coverage rises from 29.4% to 41.8%. For instance, coverage on Minerva Math improves by roughly 20%. This stronger OOD performance aligns with UVM’s design objective—capturing uncertainty to better handle unseen data.
Precision outcomes with UVM under majority voting
UVM-guided beam search does not guarantee higher precision under majority voting. In the ID settings, the average precision across models and benchmarks decreases by 1.3%, whereas in the OOD settings it increases by 3.5%. Although UVM improves coverage, these gains translate less directly to precision, leaving a substantial gap. This is consistent with intuition: because UVM explicitly models uncertainty, it naturally explores a wider range of solution paths than OVM, which can reduce the likelihood that correct paths form a majority under voting. Even so, we observe positive gains in the OOD setting, demonstrating the effectiveness of our method on unseen data.
Table 1: Performance of UVM-guided search relative to OVM-guided baseline in the ID settings (‘mst’: ‘mistral’, ‘qwm’: ‘qwen2.5-math’, ‘qw3’: ‘qwen3’)
GSM8K MATH Average
mst qwm qwm qw3
Coverage OVM 87.3% ± 0.4% 95.2% ± 0.2% 80.1% ± 0.4% 79.4% ± 0.9% 85.5%
UVM 91.6% ± 0.4% 96.9% ± 0.1% 81.4% ± 0.7% 80.5% ± 0.4% 87.6%
Precision OVM 82.2% ± 0.6% 93.6% ± 0.2% 66.5% ± 0.7% 70.9% ± 0.7% 78.3%
UVM 82.8% ± 0.3% 92.5% ± 0.3% 66.7% ± 0.8% 68.5% ± 0.8% 77.6%
Table 2: Performance of UVM-guided search relative to OVM-guided baseline in the OOD settings (Qwen2.5-Math 7B)
AIME25 Minerva Math RTN-MATH Average
Coverage OVM 8.9% ± 1.6% 23.8% ± 1.8% 55.6% ± 0.7% 29.4%
UVM 8.9% ± 1.6% 43.6% ± 1.8% 72.8% ± 0.8% 41.8%
Precision OVM 5.6% ± 1.6% 13.2% ± 1.2% 42.1% ± 0.2% 20.3%
UVM 2.2% ± 3.2% 16.5% ± 2.2% 52.7% ± 1.0% 23.8%
These results highlight the effectiveness of our method in addressing verifier failure, especially in the OOD scenarios.
6 Analysis
This section presents ablation studies on the design choices of uncertainty-aware selection and studies the effectiveness of UVM across various inference settings.
6.1 More Inference Settings
In this section, we analyze the effectiveness of our method under more inference settings in Table 3. Similar to Yu et al. (2025), we fix the number of generated paths per beam $K/b$ at 8 and conduct experiments with beam sizes of 1, 2, and 16. Under this setup, we obtain two small configurations (1 and 2) and two large configurations (16 and 32).
Table 3: Increased average coverage of UVM-guided search over OVM-guided search (‘mst’: ‘mistral’, ‘qwm’: ‘qwen2.5-math’, ‘qw3’: ‘qwen3’)
Beam Size GSM8K MATH RTN-MATH
mst qwm qwm qw3 qwm
1 -7.6% +0.4% -0.6% -10.3% -0.2%
2 -2.2% +0% -2.2% -6.2% +1.1%
16 +4.7% +0.6% +0% +0.8% +15.5%
32 +4.3% +1.7% +1.3& +1.1% +17.2%
Effectiveness across configurations UVM-guided search favors larger configurations. As shown in Table 3, UVM-guided search outperforms OVM-guided search at beam sizes of 16 and 32–for example, it yields a 4.7% gain on GSM8K using Mistral and a 15.5% gain on RTN-MATH using Qwen2.5-Math, with a beam size of 16. In contrast, the gains are limited or even negative at beam sizes of 1 and 2; for instance, on GSM8K with Mistral the gains are -7.6% and -2.2%, respectively. This pattern aligns with intuition: UVM explicitly models uncertainty and sometimes promotes high‑uncertainty candidates; while these can contain correct reasoning paths, this is not guaranteed. Consequently, larger configurations provide the necessary room for UVM’s uncertainty‑aware selection to realize its advantages.
Effectiveness under distribution shift UVM contributes more under distribution shift. As shown in Table 3, the gains rise from 0% and 1.3% with an ID training set to 15.5% and 17.2% at beam sizes of 16 and 32 when the model is trained on the RTN dataset. This highlights the effectiveness of UVM in handling uncertainty when the training set is a severe mismatch with the test set.
@@ 6.2 Ablation Study
In this section, we conduct ablation studies on the methods of uncertainty-aware selection and demonstrate the superiority of the Group Thompson Sampling algorithm in Table 4.
@@
Top-1 probability is better for candidate evaluation. As shown in Table 4, the top-1 probability ranking surpasses the UCB, demonstrating the effectiveness of using top-1 probability for candidate evaluation.
@@
Group Thompson Sampling is an effective algorithm for uncertainty-aware selection. As illustrated in Table 4, our proposed Group Thompson Sampling algorithm achieves higher coverage than the naive top-1 probability ranking. Furthermore, by sampling candidates according to the underlying top-1 probability distribution, Group Thompson Sampling eliminates the need for explicit top-1 probability estimation, making it more efficient.
These results show the superiority of our Group Thompson Sampling algorithm: It achieves higher coverage than the other baselines while requiring fewer computations, enhancing both effectiveness and efficiency compared to other uncertainty-aware selection strategies.
Table 4: Ablation on design choices of uncertainty-aware selection under coverage
GSM8K MATH Average
mst qwm qwm qw3
UCB Ranking 86.3% ± 0.4% 95.7% ± 0.3% 81.7% ± 0.7% 79.7% ± 0.3% 85.9%
Top-1 Probability Ranking 88.6% ± 0.5% 96.2% ± 0.2% 81.7% ± 0.9% 79.5% ± 1.2% 86.5%
Group Thompson Sampling 91.6% ± 0.4% 96.4% ± 0.2% 81.4% ± 0.7% 80.5% ± 0.4% 87.5%
@@ 7 Discussion
Quantifying uncertainty in search-based methods has several practical implications for LLM reasoning tasks as follows: (1) Discovering correct solutions: Correct solutions may lie along paths that are underrepresented in the training data. By incorporating uncertainty, our approach increases the likelihood of discovering these solutions. (2) Improved performance with limited additional resources: Uncertainty-aware search methods enhance performance without a significant increase in computational resources. This is because the process of quantifying uncertainty is computationally inexpensive. (3) Adaptability to real-world applications: Real-world applications often encounter OOD data. Methods that account for uncertainty are better equipped to handle such cases, enabling more reliable performance in deployment scenarios where the data may differ from training distributions.
@@ 8 Conclusion
This work addresses verifier failure in VM-guided search, a critical flaw that undermines search robustness by allowing imperfect VMs to discard promising solutions. We proposed an uncertainty-aware framework featuring UVMs to model value distributions and Group Thompson Sampling to efficiently select candidates based on this uncertainty.
Our experiments demonstrate that this framework achieves more robust performance, significantly improving solution coverage over conventional VM-guided search. The gains are especially strong on challenging OOD benchmarks like Minerva Math and our RTN-MATH setting. This confirms that modeling uncertainty is crucial for handling the distribution shift inherent in the search process.
The practical implications are significant: our method can discover correct solutions that lie along paths underrepresented in the training data, enhancing performance robustness in OOD scenarios with minimal computational overhead. While our approach boosts coverage by finding more correct answers, we note that these gains do not always translate to higher precision under majority voting, as the uncertainty-aware selection naturally promotes greater solution diversity. Future work could explore more sophisticated aggregation methods to bridge this coverage-precision gap, further harnessing the benefits of robust, uncertainty-aware search.
@@ Appendix A Appendix
$\bm{\theta}$ Parameters of the value model backbone
@@
$\bm{\zeta}$ Index vector of dimension $m$
$p_{\bm{\zeta}}$ Index distribution
@@ A.2 Detailed UVM Structure
Given the representation $\mathbf{x}$, the sampled index $\bm{\zeta}$ is mapped to the posterior value sample as:
$v=\underbrace{\mathbf{x}\mathbf{b}}_{\text{mean estimator}}+\underbrace{\mathbf{x}(u\mathbf{W}+p_{0}\mathbf{W}_{0})\bm{\zeta}^{T}}_{\text{uncertainty estimator}}$ (8)